OpenStudy (anonymous):

(square root x+7)-x=1

OpenStudy (anonymous):

\[\sqrt{x+7}-x=1\] add x to both sides \[\sqrt{x-7}=x+1\] sqare both sides \[x-7=(x+1)^2=x^2+2x+1\] set 0= 0 \[x^2+x-6=0\] factor \[(x+3)(x-2)=0\] solve \[x+3=0\] \[x=-3\] or \[x-2=0\] \[x=2\] now check

OpenStudy (anonymous):

2 works because \[\sqrt{2+7}-2=\sqrt{9}-2=3-2=1\]

OpenStudy (anonymous):

-3 does not work because \[\sqrt{-3+7}-(-3)=\sqrt{4}+3=2+3=5\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome